OpenStudy (anonymous):

Blue Chips and Red Chips A bag contains 3 blue chips and 2 red chips. Chi and Dolly play the following game. Chi selects one chip at random from the bag. If Chi selects a blue chip, Dolly gives Chi $5. If Chi selects a red chip, Chi gives Dolly $8 a)Determine Chi’s expectation. b)Determine Dolly’s expectation

ganeshie8 (ganeshie8):

for a) Chi gets money if he selects blue chip

ganeshie8 (ganeshie8):

so Chi's expectation is to draw a blue chip

OpenStudy (anonymous):

stats

OpenStudy (anonymous):

probability

OpenStudy (anonymous):

so we have five chips total

OpenStudy (anonymous):

blue is three and red is 2 so therefor we know to draw a blue chip there is 1of 3 chance and to draw a red chip there is 1 and two chance now the question is what is the probability of drawing a blue chip is 1/3

OpenStudy (anonymous):

Tell me I am on the right track here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have the form for this problem/

OpenStudy (anonymous):

there was no form just have to give the probaility

OpenStudy (anonymous):

I was wrong you have to give the expectd value E =P1·A1 + P2·A2 + P3 · A3 + … + Pn · An doe sthis help

OpenStudy (anonymous):

give me a couple min to review something here

OpenStudy (anonymous):

here is the answer: Question 436502: Blue Chips and Red Chips A bag contains 3 blue chips and 2 red chips. Chi and Dolly play the following game. Chi selects one chip at random from the bag. If Chi selects a blue chip, Dolly gives Chi $5. If Chi selects a red chip, Chi gives Dolly $8. * a) Determine Chi’s expectation. * b) Determine Dolly’s expectation. You can put this solution on YOUR website! Blue Chips and Red Chips A bag contains 3 blue chips and 2 red chips. Chi and Dolly play the following game. Chi selects one chip at random from the bag. If Chi selects a blue chip, Dolly gives Chi $5. If Chi selects a red chip, Chi gives Dolly $8. -------------------- Random gain values for Chi: 5, -8 Corresponding Probabilities: 3/5, 2/5 ------ * a) Determine Chi’s expectation. E(x) = 5(3/5)-(8)(2/5) = 3-16/5 = -1/5 = -20 cents -------------------------------------------------------- * b) Determine Dolly’s expectation. Ans: +20 cents

OpenStudy (anonymous):

hey you see the answer?